# Language Models Encode Collaborative Signals in Recommendation

### v1

@@ 4.2 Zero-shot Recommendation Performance on Entirely New Datasets (RQ2)
Task and datasets. In zero-shot recommendation [38], there is not any item or user overlap between the training set and test set [38, 33], which is different from the research line of cross-domain recommendation in ID-Rec [84]. We jointly train AlphaRec on three source datasets (i.e., Books, Movies & TV, and Video Games), while testing it on three completely new target datasets (i.e., Movielens-1M [60], Book Crossing [61], and Industrial [1]) without further training on these new datasets. (see more details about how we train AlphaRec on multiple datasets in Appendix C.3.1).
Baselines. Due to the lack of zero-shot recommenders in the field of general recommendation, we slightly modify two zero-shot methods in the sequential recommendation [85], ZESRec [37] and UniSRec [37], as baselines. We also incorporate two strategy-based CF methods, Random and Pop (see more details about these baselines in Appendix C.3.2).
@@
Industrial MovieLens-1M Book Crossing
@@
Results. Table 4 presents the zero-shot recommendation performance comparison on entirely new datasets. The best-performing methods are bold and starred, while the second-best methods are underlined. We observe that:
@@
  • AlphaRec demonstrates strong zero-shot recommendation capabilities, comparable to or even surpassing the fully trained LightGCN. On datasets from completely different platforms (e.g., MovieLens-1M and Book Crossing), AlphaRec is comparable with the fully trained LightGCN. On the same Amazon platform dataset, Industrial, AlphaRec even surpasses LightGCN, for which we attribute to the possibility that AlphaRec implicitly learns unique user behavioral patterns on the Amazon platform [1]. Conversely, ZESRec and UniSRec exhibit a marked performance decrement compared with AlphaRec. We attribute this phenomenon to two aspects. On the one hand, BERT-style LMs [4, 5] used in these works may not have effectively encoded user preference similarities, which is consistent with our previous findings in Section 2. On the other hand, components designed for the next item prediction task in sequential recommendation [86] may not be suitable for capturing the general preferences of users in CF scenarios.
@@
  • The zero-shot recommendation capability of AlphaRec generally benefits from an increased amount of training data, without harming the performance on source datasets. As illustrated in Figure 8, the zero-shot performance of AlphaRec, when trained on a mixed dataset, is generally superior to training on one single dataset [37]. Additionally, we also note that training data with themes similar to the target domain contributes more to the zero-shot performance. For instance, the zero-shot capability on MovieLens-1M may primarily stem from Movies & TV. Furthermore, we discover that AlphaRec, when trained jointly on multiple datasets, hardly experiences a performance decline on each source dataset. These findings further point to the general recommendation capability of a single pre-trained AlphaRec across multiple datasets. The above findings also offer a potential research path to achieve general recommendation capabilities, by incorporating more training data with more themes. See more details about these results in Appendix C.3.3.
4.3 User Intention Capture Performance (RQ3)
Motivation. We aim to investigate whether a straightforward paradigm shift enables pre-trained AlphaRec to perceive text-based user intentions and refine recommendations. Specifically, we introduce an adjustable hyperparameter $\alpha$ to combine user intentions with historical interests.
@@ 4.3 User Intention Capture Performance (RQ3)
Task and datasets. We test the user intention capture ability of AlphaRec on MovieLens-1M and Video Games. In the test set, only one target item remains for each user [87], with one intention query generated by ChatGPT [88, 40] (see the details about how to generate and check these intention queries in Appendix C.4.1). In the training stage, we follow the same procedure as illustrated in Section 2 to train AlphaRec. In the inference stage, we obtain the language representation $\mathbf{e}_{u}^{Intention}$ for each user intention query and combine it with the original user representation to get a new user representation as $\mathbf{\tilde{e}}_{u}^{(0)}=(1-\alpha)\mathbf{e}_{u}^{(0)}+\alpha\mathbf{e}_{u}^{Intention}$ [87]. This new user representation is sent into the freezed AlphaRec for recommendation. We report a relatively small $K$ = 5 for all metrics to better reflect the intention capture accuracy.
@@
Figure 3: User intention capture experiments on MovieLens-1M. (3(a)) AlphaRec refines the recommendations according to language-based user intention. (3(b)) The effect of user intention strength $\alpha$.
User intention capture results. Table 5 represents the user intention capture experiment results, compared with the baseline TEM [89]. Clearly, the introduction of user intention (w Intention) significantly refines the recommendations of the pre-trained AlphaRec (w/o Intention). Moreover, AlphaRec outperforms the baseline model TEM by a large margin, even without additional training on search tasks. We further conduct a case study on MovieLens-1M to demonstrate how AlphaRec captures the user (see more case study results in Appendix C.4.3). As shown in Figure 3(a), AlphaRec accurately captures the hidden user intention for “Godfather”, while keeping most of the recommendation results unchanged. This indicates that AlphaRec captures the user intention and historical interests simultaneously.
Effect of the intention strength $\alpha$. By controlling the value of $\alpha$, AlphaRec can provide better recommendation results, with a balance between user historical interests and user intent capture. Figure 3(b) depicts the effect of $\alpha$. Initially, as $\alpha$ increases, the recommendation performance rises accordingly, indicating that incorporating user intention enables AlphaRec to provide better recommendation results. However, as the $\alpha$ approaches 1, the recommendation performance starts to decrease, which suggests that the user historical interests learned by AlphaRec also play a vital role. The similar effect of $\alpha$ on Video Games is discussed in Appendix C.4.4.
@@ 5 Limitations
There are several limitations not addressed in this paper. On the one hand, although we have demonstrated the excellence of AlphaRec for multiple tasks on various offline datasets, the effectiveness of online employment remains unclear. On the other hand, although we have successfully explored the potential of language-representation-based recommenders by incorporating essential components in leading CF models, we do not elaboratively focus on designing new components for CF models.
@@ 6 Conclusion
In this paper, we explored what knowledge about recommendations has been encoded in the LM representation space. Specifically, we found that the advanced LMs representation space exhibits a homomorphic relationship with an effective recommendation space. Based on this finding, we developed a simple yet effective CF model called AlphaRec, which exhibits good recommendation performance with zero-shot recommendation and user intent capture ability. We pointed out that AlphaRec follows a new recommendation paradigm, language-representation-based recommendation, which uses language representations from LMs to represent users and items and completely abandons ID-based embeddings. We believed that AlphaRec is an important stepping stone towards building general recommenders in the future. [^1] Thc broader impact of AlphaRec will be detailed in Appendix E
@@ Appendix A Related Works
Representations in LMs. The impressive capabilities demonstrated by LMs across various tasks raise a wide concern about what they have learned in the representation space. An important and effective approach for interpreting and analyzing representations of LMs is linear probing [68, 67]. The main idea of linear probing is simple: training linear classifiers to predict some specific attributes or concepts (e.g., lexical structure [9] ) from the representations in the hidden layers of LMs. A high probing result (e.g., classification accuracy on the out-of-sample test set) tends to imply relevant information has been implicitly encoded in the representation space of LMs, although this does not imply LMs directly use these representations [68, 10]. Recent studies empirically demonstrate that concepts such as color [90], game states [8]. and geographic position are encoded in LMs. Furthermore, these concepts may even be linearly encoded in the representation space of LMs [8, 91].
@@
Collaborative filtering. Collaborative filtering (CF) [92] is an advanced technique in modern recommender systems. The prevailing CF methods tend to adopt an ID-based paradigm, where users and items are typically represented as one-hot vectors, with an embedding table used for lookup [55]. Usually, these embedding parameters are learned by optimizing specific loss functions to reconstruct the history interaction pattern [71]. Recent advances in CF mainly benefit from two aspects, graph convolution [75] and contrastive learning [92]. These CF models exhibit superior recommendation performance by conducting the embedding propagation [52, 56] and applying contrastive learning objectives [83, 63, 70]. However, although effective, these methods are still limited, due to the ID-based paradigm. Since one-hot vectors contain no feature information beyond being identifiers, it is challenging to transfer pre-trained ID embeddings to other domains [37] or to leverage leading techniques from computer vision (CV) and natural language processing (NLP) [34].
@@
LMs for recommendation. The remarkable language understanding and reasoning ability shown by LMs has attracted extensive attention in the field of recommendation. The application of LMs in recommendation can be categorized into three main approaches: LM-enhanced recommendation, LM as the modality encoder, and LLM-based recommendation. The first research direction, LLM-enhanced recommendation, focuses on empowering traditional recommenders with the semantic representations from LMs [48, 47, 46, 49, 93, 94]. Specifically, these methods introduce representations from LMs as additional features for traditional ID-based recommenders, to capture complicated user preferences. The second research line lies in adopting the LM as the text modality encoder, which is also known as a kind of modality-based recommendation (MoRec) [34, 35]. These methods tend to train the LM as the text modality encoder together with the traditional recommender. In previous studies, BERT-style LMs are widely used as the text modality encoder. The third research line, LLM-based recommendation, directly uses LLMs as the recommender and recommends items in a text generation paradigm. Early attempts focus on adopting in-context learning (ICL) [95] and prompting pre-trained LLMs [96, 97, 98, 99]. However, such naive methods tend to yield poor performance compared to traditional models. Therefore, recent studies concentrate on fine-tuning LLMs on recommendation-related corpus [16, 15, 26, 25, 29] and align the LLMs with the representations from traditional recommenders as the additional modality [17, 20, 27, 100].
Appendix B Linear Mapping
@@ B.1 Brief of Used LMs
We briefly introduce the LMs we use for linear mapping in Section 2.
  • BERT [4] is an encoder-only language model based on the transformer architecture [3], pre-trained on text corpus with unsupervised tasks. BERT adopts bidirectional self-attention heads to learn bidirectional representations.
  • RoBERTa [5] is an enhanced version of BERT. RoBERTa preserves the architecture of BERT but improves it by training with more data and large batches, adopting dynamic masking, and removing the next sentence prediction objective.
  • Llama2-7B [6] is an open-source decoder-only LLM with 7 billion parameters. Llama2 adopts grouped-query attention, with longer context length and larger size of the pre-training corpus compared with Llama-7B [101].
  • Mistral-7B [72] is an open-source pre-trained decoder-only LLM with 7 billion parameters. Mistral 7B leverages grouped-query attention, coupled with sliding window attention for faster and lower cost inference.
  • text-embedding-ada-v2 & text-embeddings-3-large [73] are leading text embedding models released by OpenAI. These models are built upon decoder-only GPT models, pre-trained on unsupervised data at scale.
  • SFR-Embedding-Mistral [74] is a decoder-based text embedding model built upon the open-source LLM Mixtral-7B [72]. SFR-Embedding-Mistral introduces task-homogeneous batching and computes contrastive loss on “hard negatives”, which brings a better performance than the vanilla Mixtral-7B model.
B.2 Extracting Representations from LMs
We present how to extract representations from LMs. For encoder-based LMs (e.g., BERT [4] and RoBERTa [5]), we use the representation of the last hidden state corresponding to the [CLS] token [40]. For decoder-based models (e.g., Llama-7B [6, 72], Mistral-7B, and SFR-Embedding-Mistral [74]), we use the representation in the last transformer block [3], corresponding to the last input token [10, 102, 73]. Especially, for the commercial closed-source model (e.g., text-embedding-ada-v2 and text-embeddings-3-large  [^2] https://platform.openai.com/docs/guides/embeddings [73]), we directly call the API interface to obtain representations.
B.3 Empirical Findings
We examinate whether the better performance of leading LMs comes from the more compact representation ability. We randomly shuffle item representations and conduct the same linear mapping experiment. As illustrated in Table 6, randomly shuffled representations, text-embeddings-3-large (Random), lags largely behind the vanilla linear mapping method, suggesting that the performance improvement not merely comes from the better feature encoding ability.
Table 6: Linear mapping performance of randomly shuffled item representations
Books Movies & TV Video Games
Recall NDCG HR Recall NDCG HR Recall NDCG HR
BERT 0.0226 0.0194 0.1240 0.0415 0.0399 0.2362 0.0524 0.0309 0.1245
text-embeddings-3-large (Random) 0.0200 0.0197 0.1316 0.0559 0.0528 0.3204 0.0562 0.0328 0.1351
text-embeddings-3-large 0.0735 0.0608 0.3355 0.1109 0.1023 0.5200 0.1367 0.0793 0.2928
Appendix C Experiments

### v2

@@ 5.2 Zero-shot Ability (Potential 2)
Experimental settings. In zero-shot recommendation, there is no item or user overlap between the training set and test set (Ding et al., 2021; Zhang et al., 2024b), which is different from the research line of cross-domain recommendation (Zhu et al., 2021). We jointly train AlphaRec on three source datasets (i.e., Books, Movies & TV, and Video Games), while testing it on three completely new target datasets (i.e., Movielens-1M (Harper & Konstan, 2016), Book Crossing (Lee et al., 2019), and Amazon Industrial & Scientific (Ni et al., 2019)) without further training on these new datasets. (see more details about training on multiple datasets in Appendix D.2.1). Due to the lack of zero-shot recommenders in general CF, we slightly modify the two zero-shot methods in the sequential recommendation (Wang et al., 2019a), ZESRec (Hou et al., 2022) and UniSRec (Hou et al., 2022), as baselines. We also incorporate two strategy-based CF methods (i.e., Random and Pop) and one method using the large language model (LLM) as zero-shot recommender (i.e., LLMRank (Hou et al., 2024b)) (see more details about baselines in Appendix D.2.2).
@@
Industrial & Scientific MovieLens-1M Book Crossing
@@
Findings. Table 4 presents the zero-shot recommendation performance comparison. The best methods are bold and starred, while the second-best methods are underlined. We observe that:
@@
  • Advanced language representations provide opportunities for learning transferable item representations. AlphaRec demonstrates strong zero-shot recommendation capabilities, comparable to or even surpassing the fully trained LightGCN. AlphaRec performs better on the Amazon Industrial & Scientific dataset, possibly because it captures user behavioral patterns of the same platform (Ni et al., 2019) through training on multiple Amazon datasets. Conversely, ZESRec and UniSRec exhibit a marked performance decrement compared with AlphaRec. We attribute this phenomenon to two aspects. On the one hand, BERT-style LMs (Devlin et al., 2019; Liu et al., 2019) used in these works may not have effectively encoded user preference similarities, which is consistent with our previous findings in Section 3. On the other hand, components designed for the next item prediction task in sequential recommendation (Kang & McAuley, 2018) may not be suitable for capturing the general preferences of users in CF scenarios. Moreover, AlphaRec also outperforms the leading LLM-based zero-shot recommender LLMRank (see Appendix D.2.3).
@@
  • The zero-shot recommendation capability of advanced language representations generally benefits from an increased amount of training data, without compromising performance on source datasets. As illustrated in Table 9, the zero-shot performance of AlphaRec, when trained on a mixed dataset, is generally superior to training on one single dataset (Hou et al., 2022). Moreover, we discover that AlphaRec, when trained jointly on multiple datasets, hardly experiences a performance decline on each source dataset. These results indicate the general recommendation capability of a single pre-trained AlphaRec across multiple datasets. The above findings also offer a potential research path to achieve general recommendation capabilities, by incorporating more training data with more themes. See more details about these results in Appendix D.2.4.
5.3 Intention-aware Ability (Potential 3)
The language understanding ability in advanced language representations (especially representations from LLM-based text embedding models) offers the opportunity for perceiving text-based user intentions and refining recommendations. To study the potential of intention-aware ability, we introduce a new hyperparameter $\alpha$ in AlphaRec to combine user intentions with historical interests.
@@ 5.3 Intention-aware Ability (Potential 3)
Experimental settings. To endow AlphaRec with user intention-aware ability, we adopt a simple paradigm shift by introducing a user intention representation $\mathbf{e}_{u}^{Intention}$. In the inference stage, we obtain the language representation $\mathbf{e}_{u}^{Intention}$ for each user intention query and combine it with the original user representation to get a new user representation as $\mathbf{\tilde{e}}_{u}^{(0)}=(1-\alpha)\mathbf{e}_{u}^{(0)}+\alpha\mathbf{e}_{u}^{Intention}$ (Ai et al., 2017). This new user representation $\mathbf{\tilde{e}}_{u}^{(0)}$ is sent into the pre-trained AlphaRec for recommendation. We test the user intention capture ability of AlphaRec on MovieLens-1M and Video Games. In the test set, only one target item remains for each user (Ai et al., 2017), with one intention query generated by ChatGPT (OpenAI, 2023; Hou et al., 2024a) (see the details about how to generate and check these intention queries in Appendix D.3.1). We report a relatively small $K$ = 5 for all metrics to better reflect the intention capture accuracy.
Findings: We report the user intention capture experiment results in Table 5, show one case study in Figure 4(a), and study the effect of $\alpha$ in Figure 4(b). We find out that:
The language understanding ability in advanced language representations enables recommenders to perceive user intentions and refine recommendations. As shown in Table 5, the introduction of user intention (w Intention) significantly refines the recommendations of the pre-trained AlphaRec (w/o Intention). Moreover, AlphaRec outperforms the baseline model TEM (Bi et al., 2020) by a large margin, even without additional training on search tasks. We further conduct a case study on MovieLens-1M to demonstrate how AlphaRec captures the user intention (see more case study results in Appendix D.3.3). Additionally, the intention-aware ability also benefits from user historical interests. Figure 4(b) depicts the effect of $\alpha$. The $\alpha$ controls the strength of user intention, where $\alpha=0$ denotes that the user intention is neglected and $\alpha=1$ denotes that the user historical interest is ignored. The convex curve in Figure 4(b) suggests that both user historical interests and user intention play vital roles. The above findings suggest the potential of adopting advanced language representations to perceive text-based user intentions and refining recommendations (see more details in Appendix D.3).
@@
Figure 4: User intention capture experiments on MovieLens-1M. (4(a)) AlphaRec refines the recommendations according to language-based user intention. (4(b)) The effect of user intention strength $\alpha$.
@@ 6 Limitations
There are several limitations unaddressed in this paper. On the one hand, while we explore the relationship between language spaces and behavior spaces through empirical results, there is no theoretical guarantee. On the other hand, although we investigate the potential of advanced language representations for recommendation, we do not design any new components for CF models.
@@ 7 Conclusion
In this paper, we explored the relationship between language space and behavior spaces for recommendation, and explored the potential for using language representations for recommendation. Empirical results suggest the possible presence of homomorphism between advanced LMs representation spaces and an effective item representation space for recommendation. Inspired by this finding, we discussed how to unleash the power of advanced language representations by developing a simple yet effective CF model called AlphaRec. Moreover, by systematically analyzing AlphaRec, we explored the potentials of advanced language representations: fast convergence, zero-shot ability, and intention-aware ability. Possible future work will involve exploring the space relationship from both theoretical and multimodal perspectives. We believed that this paper sheds light on rethinking the connection between language modeling and user behavior modeling, benefiting both natural language processing and recommender system communities.
@@ Appendix A Related Works
Representations in LMs. The impressive capabilities demonstrated by LMs across various tasks raise a wide concern about what they have learned in the representation space. Linear methods (e.g., linear probing (Merullo et al., 2023) and linear mapping (Alain & Bengio, 2017)) are important and effective approaches for interpreting and analyzing representations of LMs (Ravichander et al., 2021). The main idea of linear methods is simple: training linear classifiers to predict some specific attributes or concepts (e.g., lexical structure (Vulic et al., 2020) ) from the representations in the hidden layers of LMs, or transforming language representations into another feature space with a linear matrix. A high result of linear methods (e.g., classification accuracy on the out-of-sample test set) tends to imply relevant information has been implicitly encoded in the representation space of LMs, although this does not imply LMs directly use these representations (Ravichander et al., 2021; Gurnee & Tegmark, 2023). Recent studies empirically demonstrate that concepts such as color (Patel & Pavlick, 2022), game states (Li et al., 2023a). and geographic position are encoded in LMs. Furthermore, these concepts may even be linearly encoded in the representation space of LMs (Li et al., 2023a; Park et al., 2023).
@@
Collaborative filtering. Collaborative filtering (CF) (Ren et al., 2024a) is an advanced technique in modern recommender systems. The prevailing CF methods tend to adopt an ID-based paradigm, where users and items are typically represented as one-hot vectors, with an embedding table used for lookup (Koren et al., 2009). Usually, these embedding parameters are learned by optimizing specific loss functions to reconstruct the history interaction pattern (Rendle et al., 2012). Recent advances in CF mainly benefit from two aspects, graph convolution (Wu et al., 2023b) and contrastive learning (Ren et al., 2024a). These CF models exhibit superior recommendation performance by conducting the embedding propagation (Wang et al., 2019b; He et al., 2021) and applying contrastive learning objectives (Wu et al., 2021; Cai et al., 2023; Yu et al., 2024). However, although effective, these methods are still limited, due to the ID-based paradigm. Since one-hot vectors contain no feature information beyond being identifiers, it is challenging to transfer pre-trained ID embeddings to other domains (Hou et al., 2022) or to leverage leading techniques from computer vision (CV) and natural language processing (NLP) (Yuan et al., 2023).
@@
LMs for recommendation. The remarkable language understanding and reasoning ability shown by LMs has attracted extensive attention in the field of recommendation. The application of LMs in recommendation can be categorized into three main approaches: LM-enhanced recommendation, LM as the modality encoder, and LM-as recommender. The first research direction, LM-enhanced recommendation, focuses on empowering traditional recommenders with the semantic representations from LMs (Xi et al., 2023; Ren et al., 2024b; Wei et al., 2024; Geng et al., 2024; Chen, 2023; Wang et al., 2024; Hou et al., 2023; Mao et al., 2023; Qiu et al., 2021; Zhang et al., 2024c). Specifically, these methods introduce representations from LMs as additional features for traditional ID-based recommenders, to capture complicated user preferences. The second research line lies in adopting the LM as the text modality encoder, which is also known as a kind of modality-based recommendation (MoRec) (Yuan et al., 2023; Li et al., 2023c). These methods tend to train the LM as the text modality encoder together with the traditional recommender. In previous studies, BERT-style LMs are widely used as the text modality encoder. The third research line fails in directly using LMs as the recommender and recommends items in a text generation paradigm. Early attempts focus on adopting in-context learning (ICL) (Dong et al., 2022) and prompting pre-trained LMs (Hou et al., 2024b; Liu et al., 2023a; Dai et al., 2023; Gao et al., 2023). However, such naive methods tend to yield poor performance compared to traditional models. Therefore, recent studies concentrate on fine-tuning LMs on recommendation-related corpus (Bao et al., 2023; Zhang et al., 2023d; Lin et al., 2023b; Cui et al., 2022; Liu et al., 2024b; Hua et al., 2023; Chen et al., 2024) and align the LMs with the representations from traditional recommenders as the additional modality (Liao et al., 2024; Zhang et al., 2023e; Yang et al., 2023; Li et al., 2023d; Kong et al., 2024).
Appendix B Uncovering Collaborative Signals in LMs via Linear Mapping
@@ B.1 Brief of Used LMs
We briefly introduce the LMs we use for linear mapping in Section 3.1.
  • BERT (Devlin et al., 2019) is an encoder-only language model based on the transformer architecture (Vaswani et al., 2017), pre-trained on text corpus with unsupervised tasks. BERT adopts bidirectional self-attention heads to learn bidirectional representations.
  • RoBERTa (Liu et al., 2019) is an enhanced version of BERT. RoBERTa preserves the architecture of BERT but improves it by training with more data and large batches, adopting dynamic masking, and removing the next sentence prediction objective.
  • Llama2-7B (Touvron et al., 2023b) is an open-source decoder-only LLM with 7 billion parameters. Llama2 adopts grouped-query attention, with longer context length and larger size of the pre-training corpus compared with Llama-7B (Touvron et al., 2023a).
  • Mistral-7B (Jiang et al., 2023) is an open-source pre-trained decoder-only LLM with 7 billion parameters. Mistral 7B leverages grouped-query attention, coupled with sliding window attention for faster and lower cost inference.
  • text-embedding-ada-v2 & text-embeddings-3-large (Neelakantan et al., 2022) are leading text embedding models released by OpenAI. These models are built upon decoder-only GPT models, pre-trained on unsupervised data at scale.
  • SFR-Embedding-Mistral (Meng et al., 2024) is a decoder-based text embedding model built upon the open-source LLM Mixtral-7B (Jiang et al., 2023). SFR-Embedding-Mistral introduces task-homogeneous batching and computes contrastive loss on “hard negatives”, which brings a better performance than the vanilla Mixtral-7B model.
B.2 Generating Item Representations from LMs
We present how to extract representations from LMs. For encoder-based LMs (e.g., BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019)), we use the representation of the last hidden state corresponding to the [CLS] token (Hou et al., 2024a). For decoder-based models (e.g., Llama-7B (Touvron et al., 2023b; Jiang et al., 2023), Mistral-7B, and SFR-Embedding-Mistral (Meng et al., 2024)), we use the representation in the last transformer block (Vaswani et al., 2017), corresponding to the last input token (Gurnee & Tegmark, 2023; Todd et al., 2023; Neelakantan et al., 2022). Especially, for the commercial closed-source model (e.g., text-embedding-ada-v2 and text-embeddings-3-large  [^2] https://platform.openai.com/docs/guides/embeddings (Neelakantan et al., 2022)), we directly call the API interface to obtain representations.
Appendix C Leveraging Language Representations for Better Recommendation